# Stepwise Guided Policy Optimization: Coloring your Incorrect Reasoning in GRPO

### v4

@@ 4.4 Discussions
We highlight the motivation for evaluating both offline and online RL. In the offline setup, training uses only negative samples, allowing us to directly test whether incorrect or incomplete reasoning trajectories can improve performance. In the online setup, we simulate realistic GRPO training, where batches contain a random mix of positive and negative samples. This demonstrates that negative samples are not only effective in isolation but also remain valuable in practical settings with noisier, mixed data. While mixing positives and negatives introduces noise, simply discarding negative samples does not stabilize training; in several cases, the performance of GRPO even drops below baseline, as the model overfits to problems it can already solve.
This instability arises from limited out-of-domain generalization and catastrophic forgetting. Without exposure to challenging or partially correct reasoning, the model risks overfitting to easy cases, reinforcing shallow heuristics instead of developing robust problem-solving skills. The absence of diverse failure cases can also cause catastrophic forgetting, degrading performance on previously solvable tasks. Incorporating negative samples mitigates these issues, as SGPO consistently outperforms GRPO on Chinese OOD math benchmarks. Nonetheless, further work is needed to design more stable training frameworks, including richer reward diversification mechanisms for handling negative samples and corresponding efficient RL methods beyond GRPO.
@@ 5 Conclusion
We propose a simple and efficient framework that introduces response diversity within all-negative-sample groups and prove, in a simplified setting, that such diversification can accelerate the learning dynamic of GRPO. Empirically, our approach can yield consistent improvements across model sizes in both offline and online training over nine benchmarks, including base and distilled variants. Future works include extending theoretical results to broader multi-step reasoning tasks, applying response diversity to accelerate other RL methods, and designing lightweight, task-specific reward models that evaluate reasoning steps correctly even if they cannot solve the full problem.
@@ Acknowledgments
We sincerely appreciate Buzz High Performance Computing (https://www.buzzhpc.ai/, info@buzzhpc.ai) for providing computational resources and support for this work.
@@ Appendix A Related Works
We comment on all related topics, including reasoning through test-time compute, chain-of-thought and its variants, direct preference alignment methods, reward models and reinforcement learning from AI feedback. For an overview of more reasoning models and methods, we refer to two recent surveys [Huang and Chang, 2023, Chen et al., 2025c].
Reasoning through test-time compute.
OpenAI-o1 [Jaech et al., 2024] is among the first large-scale applications of RL to reasoning, and achieved state-of-the-art performance upon release. Following this trend, DeepSeek-R1 [Guo et al., 2025a] is the first open-weight model to match or exceed OpenAI-o1. Their real-world success stories have involved several simple yet novel techniques that enhance LLM reasoning through more test-time compute, including chain-of-thought [Wei et al., 2022], self-consistency [Wang et al., 2023], best-of-$N$ sampling [Snell et al., 2025], process reward models [Lightman et al., 2024], Monte Carlo tree search [Silver et al., 2016, Hao et al., 2023], tree-of-thought [Yao et al., 2023], and recent works on preventing overthinking [Chen et al., 2024b, Team et al., 2025, Luo et al., 2025a, Arora and Zanette, 2025] and compressing chain-of-thought [Hao et al., 2024b, Cheng and Van Durme, 2024]. More specifically, chain-of-thought is a reasoning approach where intermediate steps are explicitly written to make complex problem-solving processes more transparent and logical. Self-consistency suggests generating multiple final answers and returning the mode of an empirical distribution, enhancing test-time performance when test-time verifiers are unavailable. Unfortunately, it is computationally expensive and effective only when answers can be clustered. Best-of-$N$ sampling resolves this issue by sampling answers from the model and selecting the best at test time according to the scoring function; however, it is sensitive to the accuracy of test-time scoring functions [Gao et al., 2023]. Process reward models offer fine-grained supervision of chain-of-thought reasoning, but they might be vulnerable to reward hacking and introduce computation overhead. Monte Carlo tree search is a generic technique that allocates computational resources toward the most promising regions of the search space, and tree-of-thought and its extension [Besta et al., 2024, Gandhi et al., 2024] simplified this idea by exploring multiple reasoning paths in a specific structure, allowing language models to select the most promising line of thought for complex problem-solving. Both length regularization and compressed chain-of-thought are developed to reduce inference costs for reasoning, which is crucial for the economic feasibility, user experience and environmental sustainability of LLMs. In addition, several works have focused on specific reasoning tasks [Lampinen et al., 2024, Yang et al., 2025, Srivastava et al., 2024, Huang et al., 2025, 2024, Guo et al., 2025b, Gou et al., 2024, Wang et al., 2025], demonstrating promising performance. The recent findings Xiong et al. [2025a] have shown that the REINFORCE-type methods (including GRPO [Shao et al., 2024]) can not effectively learn from all-negative-sample groups. Our work alleviates this issue by leveraging AI feedback to differentiate negative samples. We also provide a theoretical analysis through a stylized model, explaining why such diversification improves GRPO’s learning dynamics.
@@
These methods (e.g., DPO [Rafailov et al., 2023]) are simple and stable offline alternatives to online RLHF. Various DPO variants with other objectives have been proposed, including ranking ones beyond pairwise preference data [Dong et al., 2023, Yuan et al., 2023a, Song et al., 2024, Chen et al., 2024a, Liu et al., 2025a] and simple ones that do not rely on a reference model [Hong et al., 2024, Meng et al., 2024]. Since DPO does not train a reward model, the limited size of human labels becomes a bottleneck. To alleviate this limitation, subsequent works proposed to augment preference data using a trained SFT policy [Zhao et al., 2023] or a refined SFT policy with rejection sampling [Liu et al., 2024a]. The DPO loss was recently extended to token-level MDP [Rafailov et al., 2024] given that the transition is deterministic – which has covered the fine-tuning of LLMs – and more general RL problems [Azar et al., 2024]. There are other DPO variants [Ethayarajh et al., 2024, Park et al., 2024, Xu et al., 2024, Tang et al., 2024, Meng et al., 2024, Chen et al., 2025a, Zhao et al., 2025]. For example, [Ethayarajh et al., 2024] designed the specific loss using a prospect theory, [Tang et al., 2024] optimized a general preference loss instead of the log-likelihood loss, and [Meng et al., 2024] aligned the reward function in the preference optimization objective with the generation metric. Dong et al. [2024] and [Xiong et al., 2024] proposed to generate human feedback in an online fashion to mitigate the distribution-shift and over-parameterization phenomenon. This improves DPO for complex reasoning tasks [Pang et al., 2024]. Several other works focus on unintentional alignment of DPO and developing new methods [Pal et al., 2024, Tajwar et al., 2024, Liu et al., 2024b, Xiao et al., 2024, Yuan et al., 2025, Razin et al., 2025, Chen et al., 2025b]. Among these works, [Razin et al., 2025] proposed to measure the similarity between preferred and dispreferred responses using the centered hidden embedding similarity (CHES) score and showed that filtering out preference pairs with small CHES score improves DPO, while [Chen et al., 2025b] proposed to use comparison oracles, and showed that combining it with DPO effectively alleviated the issue of unintentional alignment.
@@
For the prompt $\mathbf{x}$ with a ground-truth response $\mathbf{y}_{\mathbf{x}}^{\star}$, we evaluate by implementing a regular expression match on the final answer [Hendrycks et al., 2021]: $r(\mathbf{x},\mathbf{y})=1$ if $\mathbf{y}$ matches $\mathbf{y}_{\mathbf{x}}^{\star}$ on the final answer and $r(\mathbf{x},\mathbf{y})=0$ otherwise. An outcome reward model (ORM) [Cobbe et al., 2021, Uesato et al., 2022] is trained for estimating $r(\mathbf{x},\mathbf{y})$. In particular, we first choose $\mathbf{x}\in\mathcal{D}$ and collect training samples $(\mathbf{x},\mathbf{y}\sim\pi_{\theta}(\cdot|\mathbf{x}),r(\mathbf{x},\mathbf{y}))$. Then, we take $(\mathbf{x},\mathbf{y})$ as input and train an ORM to predict $r(\mathbf{x},\mathbf{y})$. This can be done using binary classification [Cobbe et al., 2021, Yu et al., 2024a], direct preference optimization [Hosseini et al., 2024] or next-token prediction [Zhang et al., 2024b]. Previous works also train LLMs on self-generated data using the ground-truth outcome reward model with either supervised fine-tuning [Singh et al., 2024, Yuan et al., 2023b, Zelikman et al., 2022] or online RL [Bi et al., 2024, Guo et al., 2025a]. A process reward model (PRM) is trained to score $a_{h}$ at $\mathbf{s}_{h}=(\mathbf{x},a_{1},\ldots,a_{h-1})$ either using human annotations [Lightman et al., 2024] or the value functions based on LLM-generated data [Wang et al., 2024, Luo et al., 2024, Setlur et al., 2025]; indeed, PRMs estimate either the likelihood of future success or the change in the likelihood of future success before and after taking $a_{h}$. In addition, PRMs were also developed to improve search methods [Snell et al., 2025, Wu et al., 2025], and to identify the “first pit” in an incorrect reasoning trajectory to construct preference pairs for direct preference alignment [Hwang et al., 2024, Setlur et al., 2024].

### v5

@@ 4.4 Discussions
We highlight the motivation for evaluating both offline and online RL. In the offline setup, training uses only negative samples, allowing us to directly test whether incorrect or incomplete reasoning trajectories can improve performance. In the online setup, we simulate realistic GRPO training, where batches contain a random mix of positive and negative samples. This demonstrates that negative samples are not only effective in isolation but also remain valuable in practical settings with noisier, mixed data. While mixing positives and negatives introduces noise, simply discarding negative samples does not stabilize training; in several cases, the performance of GRPO drops below baseline, as the models overfit to problems they can solve.
This instability arises from limited out-of-domain generalization and catastrophic forgetting. Without exposure to challenging or partially correct reasoning, the model risks overfitting to easy cases, reinforcing shallow heuristics instead of developing robust problem-solving skills. The absence of diverse failure cases can also cause catastrophic forgetting, degrading performance on previously solvable tasks. Incorporating negative samples mitigates these issues, and SGPO shows stronger robustness on the Chinese OOD math benchmarks we evaluate. We emphasize that SGPO does not guarantee uniform improvements on every benchmark: its gains depend on the structure of negative samples. SGPO is most beneficial when failures are due to truncated but largely correct trajectories or localized mistakes, whereas when negative samples fail very early or lack meaningful structure, the step-wise signal can be less informative and gains may diminish. These observations motivate further work on more stable training frameworks, including richer reward diversification mechanisms for handling negative samples and efficient RL methods beyond GRPO.
Benchmarking scope.
Our goal is to isolate the benefit of learning from all-negative-sample groups in outcome-based, verifiable-reward post-training with group-relative updates. We therefore benchmark SGPO primarily against GRPO under matched training pipelines (same base models, data, rollout budget, group size, KL control, and optimizer settings), since SGPO is designed as a drop-in modification of GRPO’s credit assignment only in all-negative groups. This choice makes the comparison controlled: improvements can be attributed to step-wise differentiation of negative samples rather than to changes in the broader RL recipe.
A natural question is how SGPO compares to methods that rely on process reward models (PRMs) or other step-level scoring mechanisms. Such comparisons would indeed further strengthen the empirical picture, but they introduce additional moving parts (PRM training data/labels, architecture, calibration, and inference overhead) that are orthogonal to our main question: given an outcome-verifiable setting and GRPO-style updates, can we recover learning signal from all-negative groups by cheaply localizing the first error? In this work, we keep the benchmark focused on GRPO and its outcome-based setting, and view PRM-based pipelines as complementary directions for future evaluation, especially when reliable PRMs and their training recipes are available and standardized. That said, we acknowledge that broader benchmarking (e.g., PRM-based pipelines) is an important next step, but given the recency and centrality of GRPO in outcome-based RLVR, a controlled GRPO-centered comparison is sufficient to establish relevance for the target setting.
@@ 5 Conclusion
We propose a simple and efficient framework that introduces response diversity within all-negative-sample groups and prove, in a simplified setting, that such diversification can accelerate the learning dynamic of GRPO. Empirically, SGPO improves average performance and improves coverage on harder problems, with the largest benefits in early and mid-training when all-negative groups are prevalent; gains are not uniform across benchmarks and depend on the structure/informativeness of negative samples. Future works include extending theoretical results to broader multi-step reasoning tasks, applying response diversity to accelerate other RL methods, and designing lightweight, task-specific reward models that evaluate reasoning steps correctly even if they cannot solve the full problem.
@@ Acknowledgments
We sincerely appreciate Buzz High Performance Computing (https://www.buzzhpc.ai, info@buzzhpc.ai) for providing computational resources and support for this work.
@@ Appendix A Related Works
We comment on all related topics, including reasoning through chain-of-thought and its variants, test-time compute, direct preference alignment methods, reward models and reinforcement learning from AI feedback. For an overview of more reasoning models and methods, we refer to two recent surveys [Huang and Chang, 2023, Chen et al., 2025c].
@@
Reasoning through test-time compute.
OpenAI-o1 [Jaech et al., 2024] is among the first large-scale applications of RL to reasoning, and achieved state-of-the-art performance upon release. Following this trend, DeepSeek-R1 [Guo et al., 2025a] is the first open-weight model to match or exceed OpenAI-o1. Their real-world success stories have involved several simple yet novel techniques that enhance LLM reasoning through more test-time compute, including chain-of-thought [Wei et al., 2022], self-consistency [Wang et al., 2023], best-of-$N$ sampling [Snell et al., 2025], process reward models [Lightman et al., 2024], exploration-exploitation mechanism [Chen et al., 2026b], Monte Carlo tree search [Silver et al., 2016, Hao et al., 2023], tree-of-thought [Yao et al., 2023], and recent works on preventing overthinking [Chen et al., 2024b, Team et al., 2025, Luo et al., 2025a, Arora and Zanette, 2025] and compressing chain-of-thought [Hao et al., 2024b, Cheng and Van Durme, 2024]. More specifically, chain-of-thought is a reasoning approach where intermediate steps are explicitly written to make complex problem-solving processes more transparent and logical. Self-consistency suggests generating multiple final answers and returning the mode of an empirical distribution, enhancing test-time performance when test-time verifiers are unavailable. Unfortunately, it is computationally expensive and effective only when answers can be clustered. Best-of-$N$ sampling resolves this issue by sampling answers from the model and selecting the best at test time according to the scoring function; however, it is sensitive to the accuracy of test-time scoring functions [Gao et al., 2023]. Process reward models offer fine-grained supervision of chain-of-thought reasoning, but they might be vulnerable to reward hacking and introduce computation overhead. Monte Carlo tree search is a generic technique that allocates computational resources toward the most promising regions of the search space, and tree-of-thought and its extension [Besta et al., 2024, Gandhi et al., 2024] simplified this idea by exploring multiple reasoning paths in a specific structure, allowing language models to select the most promising line of thought for complex problem-solving. Both length regularization and compressed chain-of-thought are developed to reduce inference costs for reasoning, which is crucial for the economic feasibility, user experience and environmental sustainability of LLMs. In addition, several works have focused on specific reasoning tasks [Lampinen et al., 2024, Yang et al., 2025, Srivastava et al., 2024, Huang et al., 2025, 2024, Guo et al., 2025b, Gou et al., 2024, Wang et al., 2025, Guo et al., 2025c], demonstrating promising performance. The recent findings of Xiong et al. [2025a] have shown that the REINFORCE-type methods (including GRPO [Shao et al., 2024]) can not effectively learn from all-negative-sample groups. Our work alleviates this issue by leveraging AI feedback to differentiate negative samples. We also provide a theoretical analysis through a stylized model, explaining why such diversification improves GRPO’s learning dynamics.
@@
These methods (e.g., DPO [Rafailov et al., 2023]) are simple and stable offline alternatives to online RLHF. Various DPO variants with other objectives have been proposed, including ranking ones beyond pairwise preference data [Dong et al., 2023, Yuan et al., 2023a, Song et al., 2024, Chen et al., 2024a, Liu et al., 2025a] and simple ones that do not rely on a reference model [Hong et al., 2024, Meng et al., 2024]. Since DPO does not train a reward model, the limited size of human labels becomes a bottleneck. To alleviate this limitation, subsequent works proposed to augment preference data using a trained SFT policy [Zhao et al., 2023] or a refined SFT policy with rejection sampling [Liu et al., 2024a]. The DPO loss was recently rederived and extended to a token-level MDP view [Rafailov et al., 2024], where the state is the token prefix and the transition is deterministic – which has covered the fine-tuning of LLMs – and more general RL problems [Azar et al., 2024]. There are other DPO variants [Ethayarajh et al., 2024, Park et al., 2024, Xu et al., 2024, Tang et al., 2024, Meng et al., 2024, Chen et al., 2025a, Zhao et al., 2025, Chen et al., 2026a]. For example, Ethayarajh et al. [2024] designed a DPO-style loss variant using a prospect theory, Tang et al. [2024] optimized a general preference loss instead of the log-likelihood loss, and Meng et al. [2024] aligned the reward function in the preference optimization objective with the generation metric. Dong et al. [2024] and Xiong et al. [2024] proposed to generate human feedback in an online fashion to mitigate the distribution-shift and over-parameterization phenomenon. This improves DPO for complex reasoning tasks [Pang et al., 2024]. Several other works focus on unintentional alignment of DPO and developing new methods [Pal et al., 2024, Tajwar et al., 2024, Liu et al., 2024b, Xiao et al., 2024, Yuan et al., 2025, Razin et al., 2025, Chen et al., 2025b]. Among these works, Razin et al. [2025] proposed to measure the similarity between preferred and dispreferred responses using the centered hidden embedding similarity (CHES) score and showed that filtering out preference pairs with small CHES score improves DPO, while [Chen et al., 2025b] proposed to use comparison oracles, and showed that combining it with DPO effectively alleviated the issue of unintentional alignment.
@@
For the prompt $\mathbf{x}$ with a ground-truth response $\mathbf{y}_{\mathbf{x}}^{\star}$, we evaluate by implementing a regular expression match on the final answer [Hendrycks et al., 2021]: $r(\mathbf{x},\mathbf{y})=1$ if $\mathbf{y}$ matches $\mathbf{y}_{\mathbf{x}}^{\star}$ on the final answer and $r(\mathbf{x},\mathbf{y})=0$ otherwise. An outcome reward model (ORM) [Cobbe et al., 2021, Uesato et al., 2022] is trained for estimating $r(\mathbf{x},\mathbf{y})$. In particular, we first choose $\mathbf{x}\in\mathcal{D}$ and collect training samples $(\mathbf{x},\mathbf{y}\sim\pi_{\theta}(\cdot|\mathbf{x}),r(\mathbf{x},\mathbf{y}))$. Then, we take $(\mathbf{x},\mathbf{y})$ as input and train an ORM to predict $r(\mathbf{x},\mathbf{y})$. This can be done using binary classification [Cobbe et al., 2021, Yu et al., 2024a], direct preference optimization [Hosseini et al., 2024] or next-token prediction [Zhang et al., 2024c]. Previous works also train LLMs on self-generated data using the ground-truth outcome reward model with either supervised fine-tuning [Singh et al., 2024, Yuan et al., 2023b, Zelikman et al., 2022] or online RL [Bi et al., 2024, Guo et al., 2025a]. A process reward model (PRM) is trained to score $a_{h}$ at $\mathbf{s}_{h}=(\mathbf{x},a_{1},\ldots,a_{h-1})$ either using human annotations [Lightman et al., 2024] or the value functions based on LLM-generated data [Wang et al., 2024, Luo et al., 2024, Setlur et al., 2025]; indeed, PRMs estimate either the likelihood of future success or the change in the likelihood of future success before and after taking $a_{h}$. In addition, PRMs were also developed to improve search methods [Snell et al., 2025, Wu et al., 2025], and to identify the “first pit” in an incorrect reasoning trajectory to construct preference pairs for direct preference alignment [Hwang et al., 2024, Setlur et al., 2024]. Related work such as VinePPO [Kazemnejad et al., 2024] refines process-level credit assignment in PPO without explicitly training a PRM and avoids a learned value network by using Monte Carlo rollouts from intermediate prefixes; this is complementary to our GRPO-oriented setting.